Given that f"(x)=2-2/x3 and f'(1)=0,find f'(x).Given further that f(1)=8, find f(x).


1
Expert's answer
2021-05-11T14:16:05-0400

Given that "f''(x)=2-\frac{2}{x^3}" and "f'(1)=0" , let us find "f'(x):"


"f'(x)=2x+\frac{1}{x^2}+C_1"


"0=f'(1)=2+1+C_1," and hence "C_1=-3"


"f'(x)=2x+\frac{1}{x^2}-3"


Given further that "f(1)=8" , let us find "f(x):"


"f(x)=x^2-\frac{1}{x}-3x+C_2"


"8=f(1)=1-1-3+C_2"


"C_2=11"


We conclude that "f(x)=x^2-\frac{1}{x}-3x+11."